0.04x + 0.02y = 78
x + y = 2800
---------------
solve the system and you get your answers
sum at 4% = x
sum at 2% = y
1
Expert's answer
2012-08-21T09:19:52-0400
0.04x + 0.02y = 78 (1)
x + y = 2800& (2)
---------------

(2) ==>
x = 2800 - y
(1):
0.04(2800 - y) + 0.02y = 78 ==>
112 - 0.04y + 0.02y = 78 ==>
-0.02y = -34 ==>
y = 1700.
So,
x = 2800 - 1700 = 1100.